OpenStudy (anonymous):

Please help me with this Factor completely: 18x16y10 + 27x4y5 9xy(2x16y10 + 3x4y5) 9xy(2x15y9 + 3x3y4) 9x4y5(2x4y2 + 3xy) 9x4y5(2x12y5 + 3)

OpenStudy (anonymous):

greatest common factor of \(18x^{16}y^{10} \) and \(27x^4y^5 \) is ?

OpenStudy (anonymous):

both factors contain a 9 because \(18=9\times 2\) and \(27=9\times 3\) so the greatest common factor will have a 9 in it

OpenStudy (anonymous):

greatest common factor of \(x^{16}\) and \(x^4\) is \(x^4\) since \(x^{16}=x^{12}\times x^4\) similarly \(y^{10}=y^5\times y^5\) so your greatest common factor is \(9x^4y^5\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Great thank you so much

OpenStudy (anonymous):

yw